# Protein folding with an all-to-all trapped-ion quantum computer

### v1

@@ III.1 Protein folding
  • Chignolin (GYDPETGTWG[79], a synthetic peptide that folds into a stable $\beta$-hairpin structure, a common motif in real proteins, thus widely used to validate theories. We set $\theta_{\text{cutoff}}=0.006$ for the soft and $\theta_{\text{cutoff}}=0.013$ for the hard pruning.
@@ IV Results and discussion
IV.1 Overview of IonQ’s trapped-ion quantum computers
@@
IonQ Forte and IonQ Forte Enterprise integrate acousto-optic deflectors (AOD) that allow for independent steering of each laser beam to its respective ion, substantially reducing beam alignment errors across the ion chain [94, 95]. This optical architecture, combined with a robust control system that automates calibration and optimizes gate execution, has enabled the realization of larger qubit registers with enhanced gate fidelities. Consequently, IonQ Forte-generation devices establish a scalable and high-fidelity platform for quantum information processing, marking a notable progression in trapped-ion QPU technology.

### v2

@@ III.1 Protein folding
  • Chignolin (GYDPETGTWG[79, 80], a synthetic peptide that folds into a stable $\beta$-hairpin structure, a common motif in real proteins, thus widely used to validate theories. We set $\theta_{\text{cutoff}}=0.006$ for the soft and $\theta_{\text{cutoff}}=0.013$ for the hard pruning.
@@ IV Results and discussion
IV.1 Trapped-ion quantum hardware
IonQ’s quantum processing units (QPUs), specifically the IonQ Forte and IonQ Forte Enterprise systems [93], house 36 qubits each. These qubits are realized using trapped Ytterbium ions (171Yb+), with quantum information encoded in their ground state hyperfine levels. The ions are precisely sourced through laser ablation and selective ionization, then loaded into a compact, integrated vacuum package containing a surface linear Paul trap.
@@ IV.1 Trapped-ion quantum hardware
Both QPU types integrate acousto-optic deflectors (AODs) for independent laser beam steering to each ion, significantly reducing beam alignment errors [94, 95]. This advanced optical and robust control system, which automates calibration and optimizes gate execution, has allowed IonQ to build larger qubit registers with consistently high gate fidelities. Ultimately, IonQ Forte-generation devices represent a scalable, high-fidelity platform for quantum information processing, marking a significant advancement in trapped-ion QPU technology.